OpenStudy (anonymous):

need to find an equation of the line containing the given pair of points and then simplify (-3,-6) and (-8,-2)

OpenStudy (anonymous):

\[\frac{x-(-8)}{-8-(-3)}=\frac{y-(-2)}{-2-(-6)}\]\[-\frac{x+8}{5}=\frac{y+2}{4}\]5y+4x+42=0

OpenStudy (anonymous):

okay first find the M which is gradient its = -4/5 by y2-ya/x2-x1 then put value of M in the eq y=mx+c and put coordinates of 1st set x=-3 and y=-6 you will have c which is y intercept form= -42/5 so eq of line is y=-4/5*x -42/5